OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4?

OpenStudy (anonymous):

y = -1/2x – 3 y = -1/2x + 1 y = 1/2x + 4 y = 1/2x – 1

OpenStudy (anonymous):

i think y=1/2(x-4) not x+4

OpenStudy (anonymous):

idk those are all the multiple choice

OpenStudy (anonymous):

the slope of the line y=-2x+4 is -2 as the line we required is perpendicular to this line then product of slopes are -1 m1*m2=-1 m1*-2=-1 m1=1/2 then we get the slope of our required is 1/2 now the point slope form is (y-y1)=m(x-x1) now y-1=1/2(x-4) y=1/2(x-4) + 1 when we take lcm and simplify we get y=1/2(x-2) but there is no option i couldn't understand